i am working on negative exponents and i dont know how to solve -3 to the exponent of 9 dived by -3 to the power of 9 im stuck can you please help me?
Answer by stanbon(75887)

i am working on negative exponents and i dont know how to solve -3 to the exponent of 9 divided by -3 to the power of 9 im stuck can you please help me?
------------
This problem does not have a negative exponent:
------------------
(-3)^9/(-3)^9 = 1
---
(-3)^9/-3^9 = [(-1)^9*(3^9)]/[-(3^9)] = [-(3^9)/-(3^9)] = 1
=================
-3^9/-3^9 = 1
=================
Note: A negative in the exponent means you should
invert the base.
===================
Cheers,
Stan H.
============